OpenStudy (anonymous):

Consider the paraboloid z=x^2+y^2. The plane 6x−6y+z−8=0 cuts the paraboloid, its intersection being a curve. Find "the natural" parametrization of this curve.

OpenStudy (anonymous):

I have x^2 + y^2 = 6y - 6x + 8

OpenStudy (anonymous):

no clue what to do now

ganeshie8 (ganeshie8):

Yes parameterize that circle

OpenStudy (anonymous):

x^2 -6x + y^2 -6y = 8 is a circle?

OpenStudy (anonymous):

i thought a circle was x^2 + y^2 = R^2

ganeshie8 (ganeshie8):

\[(x-h)^2 + (y-k)^2 = r^2\] is a circle with radius \(r\) and center \((h, k)\)

OpenStudy (anonymous):

hmm, ok so i have to complete the square

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

AH

ganeshie8 (ganeshie8):

see if this helps https://answers.yahoo.com/question/index?qid=20080221162628AAGcxo6 i don't remember "natiral parameterization", il need to review i guess

OpenStudy (anonymous):

ah thank you :)

OpenStudy (anonymous):

I have (x-3)^2 + (y-3)^2 = 26

ganeshie8 (ganeshie8):

that looks good!

ganeshie8 (ganeshie8):

do we have \(x-3 = \sqrt{26}\cos t\) \(y-3 = \sqrt{26}\sin t\) ?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

@ganeshie8 not working :(

ganeshie8 (ganeshie8):

what are you entering?

OpenStudy (anonymous):

x(t) = 3 + sqrt(26) cos(t)

ganeshie8 (ganeshie8):

what about y and z

OpenStudy (anonymous):

y(t) = 3+sqrt(26)sin(t) worked though

OpenStudy (anonymous):

nvm got it, sign error, ugh !

OpenStudy (anonymous):

@ganeshie8 , ty :)

ganeshie8 (ganeshie8):

nice nice